OpenStudy (anonymous):

Carla purchased a hat for $12, a purse for $26, and a dress for $125. The sales tax rate was 5.6 percent. What was the total amount she paid?

OpenStudy (anonymous):

Add all the items together. To find the amount of tax for the total: multiply your total amount by the sales tax and add the number you get to your total for the items. \[12 + 26 + 125 = 163 * 0.056 (tax /100) =9.128 + 163.00 = 172.128 \approx $172.13 total. \]